the perimeter of a rectangle is 160 ft. One fourth the length is the same as twice the width. Find the dimensions of the rectangle.
Answer by Abbey(339)

perimeter = 2L+2W
160=2L+2W
Divide by 2
80=L+W
Additionally we know that
1/4L=2W
Multiply by 4
L=8W
So we can substitute that in:
80=8W+W
80=9W
8.88=W
L=8W
L=8(8.88)
L=71.77
160=2(71.11)+2(8.88)
Pretty unusual numbers for this kind of math problem...